# Random Search as a Baseline for Sparse Neural Network Architecture Search

### 13 Mar 2024 v1

@@
Sparse neural networks have shown similar or better generalization performance than their dense counterparts while having higher parameter efficiency. This has motivated a number of works to learn, induce, or search for high performing sparse networks. While reports of quality or efficiency gains are impressive, standard baselines are lacking, therefore hindering having reliable comparability and reproducibility across methods. In this work, we provide an evaluation approach and a naive Random Search baseline method for finding good sparse configurations. We apply Random Search on the node space of an overparameterized network with the goal of finding better initialized sparse sub-networks that are positioned more advantageously in the loss landscape. We record sparse network post-training performances at various levels of sparsity and compare against both their fully connected parent networks and random sparse configurations at the same sparsity levels. We observe that for this architecture search task, initialized sparse networks found by Random Search neither perform better nor converge more efficiently than their random counterparts. Thus we conclude that Random Search may be viewed as a suitable neutral baseline for sparsity search methods.
@@ 2 The Method
The Weedout training process consists of two phases, (1) the Weedout phase, and (2) the sparse network learning phase. First, a fully connected overparameterized neural network is initialized with random weights. A number of sparsely-connected sub-networks are then sampled from this fully connected network forming a population of sparse networks. Once the population is established, Random Search is applied to the population for finding the best solution among this population. We call this phase the Weedout phase. After finding the best solution, the found sparse network is then trained normally with SGD as part of a normal learning phase. The Weedout phase is further discussed below.
Genrally, any RSH method may be applied for the purpose of finding a better sparse network among the candidate population. Suitability and effectiveness of any given RSH strategy depends on a variety of factors, such as the given learning task, the population count, number of iterations for the strategy, etc. Random Search is a basic strategy and suitable as a relatively naive baseline compared to other search heuristics such as Binary Tournament which is commonly used in Evolutionary Computation applications [32].
@@
The process of creating random sparse sub-networks may be analogous to the creation of randomly connected networks such as in the work of Xie et al. [34]. However, creation of random networks from an overparameteried network creates additional opportunities to benefit from tensor-favorable hardware and software platforms without requiring computationally expensive graph computing facilities.
@@ 3 Experiments
We conduct experiments to study the application of Weedout on an image classification task. We measure both train and test classification accuracy of fully trained sub-networks found by Weedout at various levels of larger network sparsity. For control, we compare sparse networks found by Weedout versus a randomly drawn sparse network from the population. This network can be conceptually thought of as a network with "fixed dropout nodes". Several of the mentioned work [12, 13, 14, 15, 16, 17, 34] run experiments on the CIFAR-10 image classification task using a VGG-style [35] convolutional network architecture. We conduct our experiments on the CIFAR-10 task, but using the Wide ResNet architecture [36] which generally is more reproducible and exhibits higher performance on single GPU machines than the VGG architectures.
@@ 4 Results and Discussion
For these Wide ResNet architectures, we observed a progressive deterioration of prediction accuracy on the test set as sparsity increases as shown Figure 2. Interestingly, there is only approximately 15% drop in test accuracy between a fully dense network and one with 80% sparsity. This can be a reasonable trade-off for applications where parameter efficiency is of critical importance.
@@
We conclude that while Random Search is able to find networks with better initialization, this advantage is quickly overran by gradient-based learning over multiple passes on the training dataset. Hence any method that can successfully find an initial sparse configuration that performs better than a randomly initialized configuration post training is at least better than Random Search.
@@ 5 Future Work
in our experiments we have applied Weedout to produced sparsity uniformly and layer-wise in all layers of the Wide ResNet convolutional neural network. Further experimentation may yield informative results on how non-uniform sparsity affects performance, e.g. a higher level of sparsity on lower-level feature extraction layers, versus less sparsity in higher representation space layers. Particularly for convolutional neural networks, sparsity can also be extended to channels, e.g. as done in some surveyed works [14, 40].
@@
the sparsity search process can be continued through the entire training process, i.e. similar to Dropout [41], but where finding the Dropout binary mask is under search. We envision the opportunity of finding effective masks while continuing to update the network weights, and arriving at a mask which we can be immediately used to prune the overparameterized network. Of course in this case, we lose the opportunity of finding an efficient sparse sub-network prior to training and gaining training compute performance efficiency.

### 14 Mar 2024 v2

@@
Sparse neural networks have shown similar or better generalization performance than their dense counterparts while having higher parameter efficiency. This has motivated a number of works to learn or search for high performing sparse networks. While reports of task performance or efficiency gains are impressive, standard baselines are lacking leading to poor comparability and unreliable reproducibility across methods. In this work, we propose Random Search as a baseline algorithm for finding good sparse configurations and study its performance. We apply Random Search on the node space of an overparameterized network with the goal of finding better initialized sparse sub-networks that are positioned more advantageously in the loss landscape. We record the post-training performances of the found sparse networks and at various levels of sparsity, and compare against both their fully connected parent networks and random sparse configurations at the same sparsity levels. First, we demonstrate performance at different levels of sparsity and highlight that a significant level of performance can still be preserved even when the network is highly sparse. Second, we observe that for this sparse architecture search task, initialized sparse networks found by Random Search neither perform better nor converge more efficiently than their random counterparts. Thus we conclude that Random Search may be viewed as a reasonable neutral baseline for sparsity search methods.
@@ 2 The Method
The Weedout training process consists of two phases, (1) the Weedout phase, and (2) the sparse network learning phase. First, a fully connected overparameterized neural network is initialized with random weights. A number of sparsely-connected sub-networks are then sampled from this fully connected network forming a population of sparse networks. Once the population is established, Random Search is applied to the population for finding the best solution among this population. We call this phase the Weedout phase. After finding the best solution, the sparse network is then trained normally with SGD as part of a normal learning phase. The Weedout phase is further discussed below.
Generally, any RSH method may be applied for the purpose of finding a better sparse network among the candidate population. Suitability and effectiveness of any given RSH strategy depends on a variety of factors, such as the given learning task, the population count, number of iterations for the strategy, etc. Random Search is a basic strategy and suitable as a relatively naive baseline compared to other search heuristics such as Binary Tournament which is commonly used in Evolutionary Computation applications [32].
@@
The process of creating random sparse sub-networks may be analogous to the creation of randomly connected networks such as in the work of Xie et al. [34]. However, creation of random networks from an overparameterized network creates additional opportunities to benefit from tensor-favorable hardware and software platforms without requiring computationally expensive graph computing facilities.
@@ 3 Experiments
We conduct experiments to study the application of Weedout on an image classification task. We measure both train and test classification accuracy of fully trained sub-networks found by Weedout at various levels of larger network sparsity. For control, we compare sparse networks found by Weedout versus a randomly drawn sparse network from the population. This network can be conceptually thought of as a network with "fixed dropout nodes". Several of the mentioned work [12, 13, 14, 15, 16, 17, 34] run experiments on the CIFAR-10 image classification task using a VGG-style [35] convolutional network architecture. We also conduct our experiments on the CIFAR-10 task, but use the Wide ResNet architecture [36] which generally is more reproducible and exhibits higher performance on single GPU machines than the VGG architectures.
@@ 4 Results and Discussion
For these Wide ResNet architectures, we observed a progressive deterioration of prediction accuracy on the test set as sparsity increases as shown in Figure 2. Interestingly, there is only approximately 15% drop in test accuracy between a fully dense network and one with 80% sparsity. This can be a reasonable trade-off for applications where parameter efficiency is of critical importance.
@@
We conclude that while Random Search is able to find networks with better initialization, this advantage is quickly overrun by gradient-based learning over multiple passes on the training dataset. Hence any method that can successfully find an initial sparse configuration that performs better than a randomly initialized configuration post training is at least better than Random Search.
@@ 5 Future Work
in our experiments we have applied Weedout to produce sparsity uniformly and layer-wise in all layers of the Wide ResNet convolutional neural network. Further experimentation may yield informative results on how non-uniform sparsity affects performance, e.g. a higher level of sparsity on lower-level feature extraction layers, versus less sparsity in higher representation space layers. Particularly for convolutional neural networks, sparsity can also be extended to channels, e.g. as done in some surveyed works [14, 40].
@@
the sparsity search process can be continued through the entire training process, i.e. similar to Dropout [41], but where finding the Dropout binary mask is under search. We envision the opportunity of finding effective masks while continuing to update the network weights, and arriving at a mask which can be immediately used to prune the overparameterized network. Of course in this case, we lose the opportunity of finding an efficient sparse sub-network prior to training and gaining training compute performance efficiency.